OpenStudy (anonymous):

At a high temperature, the equilibrium constant for the decomposition of hydrogen iodide is 65.0. If the initial concentration of HI is 1.60 M, what is the concentration of hydrogen at equilibrium? 2Hl(g) H2(g) + I2(g)

OpenStudy (cuanchi):

It is the same that the previous problem 1.4234= 1.42 2HI ----> H2 + I2 I 1.6M 0.0 0.0 C -x +x +x E 1.6-x x x Keq=[H2][I2]/[HI]^2= x^2/(1.6-x)^2 =(x/(1.6-x)^2 calculate the value for x \[x= \frac{ 1.6\sqrt{65} }{ 1+\sqrt{65} }=1.4234\]